### v1

# EEGDM: Learning EEG Representation with Latent Diffusion Model

Abstract
While electroencephalography (EEG) signal analysis using deep learning has shown great promise, existing approaches still face significant challenges in learning generalizable representations that perform well across diverse tasks, particularly when training data is limited. Current EEG representation learning methods including EEGPT and LaBraM typically rely on simple masked reconstruction objective, which may not fully capture the rich semantic information and complex patterns inherent in EEG signals. In this paper, we propose EEGDM, a novel self-supervised EEG representation learning method based on the latent diffusion model, which leverages EEG signal generation as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. EEGDM incorporates an EEG encoder that distills EEG signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals. This design endows EEGDM with a compact latent space, which not only offers ample control over the generative process but also can be leveraged for downstream tasks. Experimental results show that EEGDM (1) can reconstruct high-quality EEG signals, (2) effectively learns robust representations, and (3) achieves competitive performance with modest pre-training data size across diverse downstream tasks, underscoring its generalizability and practical utility.
Introduction
Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes. The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al. 2020). Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe 2011).
Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Zhao, Yan, and Lu 2021; Zheng et al. 2023; Chen et al. 2025). Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models. Task-specific models face significant challenges in generalization due to substantial variations across datasets in channel configurations, sampling rates, and EEG recording time lengths. These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability. For multiple tasks, two large EEG models, EEGPT (Wang et al. 2024) and LaBraM (Jiang, Zhao, and Lu 2024) have recently been proposed. Both models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks. However, LaBraM uses a substantially larger dataset, and EEGPT’s performance across diverse specific tasks still leaves room for improvement.
In this paper, we introduce the Diffusion Model (DM) (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM. Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics. Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution. This makes DM potentially valuable for generic EEG representation learning. Building on this observation, our framework uses an EEG encoder that transforms input signals and their channel augmentations into compact latent representations, which condition the diffusion process. This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al. 2013; Cole and Voytek 2019). To address these challenges, we incorporate two key technical components in our EEGDM framework. First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information. Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers 2013). This transformation mitigates the challenge of reconstructing the original noisy signal. Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al. 2022) and Diffusion Transformers (DiTs) (Peebles and Xie 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space. This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks. Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks with modest pre-training data requirements and exhibiting strong generalizability.
Our contributions are summarized as follows:
  • We propose EEGDM, a novel diffusion model-based self-supervised framework for EEG representation learning. By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics, even with limited pre-training data.
  • EEGDM incorporates channel augmentation and elaborately designed PCA-based latent space operations to (1) enhance the conditional information available for DM, (2) enable robust EEG signal reconstruction, and (3) improve the cross-dataset generalizability of learned EEG representations.
  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (a) competitive performance compared to existing methods and (b) strong cross-dataset transferability and practical utility in EEG applications.
Related Work
Self-supervised EEG Representation Learning
Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations. This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova 2019; He et al. 2022). However, its full potential in EEG analysis remains largely unexplored (Jiang et al. 2025).
Some studies (Mohsenvand et al. 2020; Banville et al. 2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al. 2020) to time series data. Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al. 2024). For instance, BENDR (Kostas et al. 2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part. LaBraM (Jiang, Zhao, and Lu 2024) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook. EEGPT (Wang et al. 2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
Unlike the above approaches, we explore self-supervised EEG representation learning via generative models. Our proposed EEGDM not only facilitates high-quality EEG generation but also enables the learning of rich and robust representations for downstream tasks.
Diffusion Model
Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics. This fundamental characteristic endows them with dual capabilities: generating high-quality samples and learning meaningful representations inherent in the data distribution. The pioneering work by Sohl-Dickstein et al. (Sohl-Dickstein et al. 2015) and subsequent advances by Ho et al. (Ho, Jain, and Abbeel 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al. 2022; Peebles and Xie 2023; Zhang, Rao, and Agrawala 2023). More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al. 2023; Chen et al. 2024b; Hudson et al. 2024).
In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al. 2024; Chen et al. 2024a; Huang, Li, and Chen 2025), or a few specific supervised tasks (Kim et al. 2023, 2024). However, the potential of diffusion models for generic EEG representation learning remains largely unexplored. This presents a significant opportunity, as learning the complex temporal and spatial characteristics of EEG signals can benefit from the understanding of rich data distributions inherent in diffusion models’ denoising process. In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
Method
Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying properties, structures, and dynamics of the data during the generative process. Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
Figure 1: The overall architecture of EEGDM. The original EEG data is projected onto a latent space using PCA. Noise is added in this latent space, and the denoised latent representation is predicted by learning the conditional latent diffusion model. The prediction is then projected back to the original space by inverse PCA to obtain the generated target EEG signal. The encoder transforms the source EEG signal (including original data and augmentations) into a concise EEG representation. It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
Framework Architecture
Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation. Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks. It is worth noting that in the denoising process of DM, directly reconstructing raw EEG signals is challenging due to their inherently low SNR. To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR. The overall architecture of EEGDM is illustrated in Figure 1.
Specifically, given the original multi-channel EEG signals $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of EEG electrodes (channels) and $T$ is the total timestamp, we segment $X$ into ${\lfloor{\frac{T-t^{s}}{s^{t}}}\rfloor}+1$ samples. This segmentation assumes a timestamp length [^1] Here superscript is used; subscript is reserved for diffusion formation in subsequent sections. $t^{s}$ per sample and a stride $s^{t}$. As shown in Figure 1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$. Our diffusion model operates on this latent space, where noise is added and predicted. The denoised latent representation predicted by the diffusion model is projected back to the original space by inverse PCA to obtain the target EEG signal. Owing to the effectiveness of the DiT model in the latent space, we utilize it to predict the denoised latent representations. Furthermore, we introduce an EEG encoder that transforms the input source signal $x^{\prime}$ (including original EEG sample and its augmentations) into an EEG representation containing rich semantic information. The sum of this EEG representation and the noise timestep representation (i.e., timestep $t$ in the noise-adding process) encoded by the time embedding layer serves as the conditional input information for the diffusion model. This diffusion model design can utilize the semantically rich external EEG representation for the generative task, leading to a significant improvement in generative performance. Consequently, this enhances the model’s representation learning capability.
Figure 2: The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
The EEG Encoder
The representation output from the EEG encoder lies at the core of EEGDM, which (1) acts as the communication bridge between the encoder and the denoiser ($\epsilon_{\theta}$), and (2) guides the denoising process. In more details, this output represents a clean source EEG signal semantically related to the target EEG data. The introduction of the encoded EEG representation is to enable hidden states of the diffusion model to predict noise-invariant and semantically meaningful EEG representations from noisy inputs, guiding the reconstruction of the target EEG data during the denoising process.
The core motivation behind the above EEG encoder design arises from the inherently under-determined nature of the denoising task. The denoiser tends to exploit any relevant knowledge or informative cues to infer and recover the original data distribution as accurately as possible. This incentivizes the encoder to distill salient commonalities between the source and target EEG signals into a compact representation. We constrain the encoded EEG representation to a low-dimensional latent space and guide the denoiser through feature modulation. This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser. Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov 2006; Higgins et al. 2017), our strategy enables the encoder to learn a representation that supports latent space denoising. This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG. Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy 2021) that operates on the sequences of tokens. We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences. We retain many of ViT’s best practices, enabling the EEG encoder to be a Transformer architecture that can process input EEG signals with various formats (different sampling rates, number of channels, and time lengths). The input of the encoder is the source EEG signal $x^{\prime}$, which consists of the original samples and the EEG channel augmentations. EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data. Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al. 2020) to be the most effective for extracting useful features.
Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$). The total number of patches for $x^{\prime}$ is ${\left|x^{\prime}\right|}=C{\lfloor{\frac{t^{s}}{\omega}}\rfloor}$. Then a temporal encoder (i.e., a temporal convolution block) is used to encode each EEG patch into a patch embedding $e^{\prime}=\{e^{\prime}_{j,k}\in\mathbb{R}^{d}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$, where $d$ is the dimension of the embeddings. The learnable positional embeddings are added to all input patch embeddings. Finally, the sequence of embeddings is fed directly into the Transformer network. We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model. In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation. As shown in Figure 2, we train a linear classifier as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ cross-entropy loss. The implementation details can be found in the supplementary material.
EEG Data Generation by Diffusion Model
As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging. To address this, we first enhance the SNR of the raw EEG signals using PCA. Similar to the LDMs (Rombach et al. 2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals. In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process. After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
Diffusion Formulation
The input to our proposed latent diffusion model, which is conditioned on channel augmentation, is the latent representation $z=Px$ obtained through PCA pre-processing. Here, we use the $\omega$-length window to segment each $x$ before applying PCA to obtain the corresponding latent representation. The PCA basis vectors are computed via eigen-decomposition without the need for gradient-based training. Our diffusion model operates within this latent space. The noisy input $z_{t}$ at a given timestep is transformed into patches, which are then linearly embedded into a sequence of tokens with dimension $d$ (implemented by the patchify layer). Here, we apply a sine-cosine version of the positional embeddings to all input tokens. Then, the input tokens are processed by several DiT blocks. By applying the reparameterization trick, we can do the sampling as follows:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (1)
where $\epsilon_{t}\sim{\mathcal{N}(0,\mathbf{I})}$. The diffusion model is trained to learn the reverse denoising process:
$p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$ (2)
where the network is used to predict the statistics of $p_{\theta}$. By reparameterizing $\mu_{\theta}$ as a noise prediction network (i.e., denoiser) $\epsilon_{\theta}$, the model can be trained using the mean squared error between the predicted noise $\epsilon_{\theta}(z_{t})$ and the ground truth sampled Gaussian noise $\epsilon_{t}$:
$\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$ (3)
To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al. (2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$. After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction. Both outputs have the same data shape as the input. Once $p_{\theta}$ is trained, new latent representations can be sampled by initializing $z_{t_{max}}\sim{\mathcal{N}(0,\mathbf{I})}$ and sampling $z_{t-1}\sim{p_{\theta}(z_{t-1}|z_{t})}$ via the reparameterization trick. Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
$\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$. (4)
Conditional Guidance
We incorporate the encoded EEG representation as conditional input for the diffusion model. It guides the generative process and is further leveraged for downstream tasks. To this end, we broadly consider EEG signals in this context as any collection of EEG recordings that exhibit meaningful interrelationships (e.g., semantic correlations), particularly those derived through various augmentations of source EEG signals — an approach well-established in contrastive learning frameworks (Chen et al., 2020). Specifically, each augmented EEG signal is encoded using a shared EEG encoder, and the resulting EEG representations $e^{1:m}$ are aggregated into a single vector $e$ by taking their mean, enabling the model to incorporate richer conditional information. When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$. Under this setting, classifier-free guidance (Ho and Salimans 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized. According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$. By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
$\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$ (5)
$\widetilde{\epsilon}_{\theta}(z_{t},e)=\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\nabla_{z}\log p(z|e))},$
where $s>1$ denotes the scale of the guidance. The diffusion model with $e=\emptyset$ is trained by randomly dropping $e$ during training and replacing it with a learned null embedding $\emptyset$. The classifier-free guidance sampling method can significantly improve the quality of generated samples by effectively utilizing the conditional information we introduced.
Methods Model Size Balanced Accuracy ↑ AUROC ↑
[…]
Table 1: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
Methods Model Size Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
[…]
Table 2: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
Datasets Methods Balanced Accuracy ↑ Cohen’s Kappa ↑ Weighted F1 / AUROC ↑
[…]
Table 3: Performance comparison on BCIC. The best metric is highlighted in bold, and the second-best metric is underlined.
Variants PCA EEG Augmentation BCIC-2A ↑ BCIC-2B ↑ TUAB (10%) ↑ TUEV (10%) ↑
[…]
Table 4: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
Conditional Input Setup
The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models. We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$). Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$. Additionally, we regress the dimension-wise scaling parameters $\alpha$, which are applied just before each residual connection in the DiT blocks. The MLP for all $\alpha$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function. In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
Experiments
Datasets
Pre-training datasets. We pre-train the proposed EEGDM on several publicly available EEG datasets covering a range of paradigms, including PhysioMI (motor imagery and execution dataset) (Goldberger et al. 2000), TSU (steady-state visual evoked potential dataset) (Wang et al. 2016), SEED (emotion recognition dataset) (Zheng and Lu 2015), and M3CV (multi-subject multi-session multi-paradigm dataset) (Huang et al. 2022).
Datasets of downstream specific tasks. (1) TUAB (abnormal detection) (Obeid and Picone 2016): A corpus of EEG recordings annotated as either normal or abnormal. (2) TUEV (event type classification) (Obeid and Picone 2016): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG). (3) BCIC-2A (motor imagery classification) (Tangermann et al. 2012): A four-class motor imagery dataset comprising left hand, right hand, feet, and tongue movement imagination tasks. (4) BCIC-2B (motor imagery classification) (Steyrl et al. 2016): A binary motor imagery dataset containing left-hand and right-hand movement imagination tasks.
Experimental Setup
We evaluate EEGDM’s representation learning capabilities through fine-tuning on downstream tasks. Following previous work (Wang et al. 2024), we adopt standard preprocessing procedures including resampling, normalization, segmentation, and flattening. For TUAB and TUEV datasets, we use the provided train-test splits, while BCIC-2A and BCIC-2B employ Leave-One-Subject-Out (LOSO) cross-validation. The reported model sizes correspond to the downstream networks used for fine-tuning, as the generative modules are not involved in downstream tasks. We compare with state-of-the-art baselines and provide results for both EEGDM (25M parameters) and a compact variant, EEGDM-Tiny (3M parameters), which uses a smaller pre-training architecture. Detailed configurations for both model variants are provided in the supplementary material. Complete implementation details and hyperparameters for both pre-training and fine-tuning are also provided in the supplementary material.
Results and Analysis
In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis. All experiments are conducted with three random seeds, and we report mean and standard deviation values as shown in Tables 1, 2, and 3. Additional results under different fine-tuning settings are provided in the supplementary material.
On the TUAB dataset, EEGDM achieves a balanced accuracy of 80.17%, outperforming both the previous state-of-the-art method EEGPT (79.83%) and other baselines. The AUROC score of 87.39% further demonstrates strong discriminative capability. For TUEV classification, EEGDM obtains 60.81% balanced accuracy with a Cohen’s Kappa coefficient 0.5218, achieving either (1) good balance between performance and pre-training data size, when compared to EEGPT, or (2) substantial improvements over other baselines. On motor imagery classification, EEGDM shows competitive performance, achieving 55.95% balanced accuracy on BCIC-2A and 72.19% on BCIC-2B, with the latter outperforming both EEGPT and LaBraM.
Regarding the comparison with EEGPT and LaBraM, the following detailed explanations are necessary. Following EEGPT (Wang et al. 2024), all comparisons are conducted with LaBraM-Base (Jiang, Zhao, and Lu 2024). The LaBraM results are based on the same pre-training dataset as ours, rather than their original larger dataset (which includes TUEG dataset, whose distribution is similar to the TUAB and TUEV). Remarkably, our EEGDM achieves performance comparable to or better than EEGPT despite using a smaller pre-training dataset. Even our compact EEGDM-Tiny model maintains competitive performance while using substantially fewer parameters. These results validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
Ablation Study
We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table 4, in which the average balanced accuracy on three random seeds is reported. For BCIC-2A and BCIC-2B, we report the results using the LOSO evaluation protocol. Variant D consistently achieves the best performance across all settings, demonstrating that both PCA and EEG augmentation are important for our model. Furthermore, for TUAB and TUEV, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last two columns in Table 4); we also use this setting in Figure 3. The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
We also investigate the impact of the number of PCA components on downstream task performance. Figure 3 presents results on TUEV and BCIC-2B with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance. In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance. To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
To evaluate generation quality, we use the diffusion model in EEGDM under the same pre-training configuration, and compute Pearson correlation coefficients between the generated and original signals in both time and frequency domains. When PCA is applied, the correlation coefficients improve significantly to 0.36 (time domain) and 0.88 (frequency domain), compared to lower values of 0.003 and 0.83 without PCA. These results demonstrate that integrating PCA substantially improves the quality of EEG data generation, which further facilitates better representation learning.
Figure 3: Performance comparison with different numbers of PCA components on TUEV and BCIC-2B datasets.
Visualization
To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton 2008) as presented in Figure 4. We randomly sample EEG signals from the test sets of four downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder. The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns. This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks. Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses provided in the supplementary material.
(a) TUAB
(b) TUEV
(c) BCIC-2A
(d) BCIC-2B
Figure 4: The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by the visualizations in supplementary material; (2) Effective representation learning, evidenced by distinct class separation in t-SNE visualizations and consistent performance improvements across tasks; (3) Competitive performance across diverse downstream tasks using modest pre-training data sizes, particularly given that EEGDM matches or exceeds EEGPT’s performance while using smaller pre-training datasets.
Conclusion
We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation. Our EEGDM takes the generative task as the self-supervised EEG representation learning objective. By (1) integrating a denoiser with the EEG encoder and (2) making use of PCA-based projection and EEG channel augmentation, EEGDM can perform the generation task in the latent space using rich conditional information while learning powerful EEG semantic representations that can be applied to diverse downstream tasks. Comprehensive experimental results demonstrate the model’s effectiveness across multiple downstream tasks.

### v2

@@
 Abstract
-While electroencephalography (EEG) signal analysis using deep learning has shown great promise, existing approaches still face significant challenges in learning generalizable representations that perform well across diverse tasks, particularly when training data is limited.
-Current EEG representation learning methods including EEGPT and LaBraM typically rely on simple masked reconstruction objective, which may not fully capture the rich semantic information and complex patterns inherent in EEG signals.
-In this paper, we propose EEGDM, a novel self-supervised EEG representation learning method based on the latent diffusion model, which leverages EEG signal generation as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics.
-EEGDM incorporates an EEG encoder that distills EEG signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals.
+Recent advances in self-supervised learning for EEG representation have largely relied on masked reconstruction, where models are trained to recover randomly masked signal segments.
+While effective at modeling local dependencies, such objectives are inherently limited in capturing the global dynamics and long-range dependencies essential for characterizing neural activity.
+To address this limitation, we propose EEGDM, a novel self-supervised framework that leverages latent diffusion models to generate EEG signals as an objective.
+Unlike masked reconstruction, diffusion-based generation progressively denoises signals from noise to realism, compelling the model to capture holistic temporal patterns and cross-channel relationships.
+Specifically, EEGDM incorporates an EEG encoder that distills raw signals and their channel augmentations into a compact representation, acting as conditional information to guide the diffusion model for generating EEG signals.
 This design endows EEGDM with a compact latent space, which not only offers ample control over the generative process but also can be leveraged for downstream tasks.
-Experimental results show that EEGDM (1) can reconstruct high-quality EEG signals, (2) effectively learns robust representations, and (3) achieves competitive performance with modest pre-training data size across diverse downstream tasks, underscoring its generalizability and practical utility.
-Introduction
+Experimental results show that EEGDM (1) reconstructs high-quality EEG signals, (2) learns robust representations, and (3) achieves competitive performance across diverse downstream tasks, thus exploring a new direction for self-supervised EEG representation learning.
+Index Terms:
+Self-supervised learning, diffusion model, EEG representation learning.
+I Introduction
 Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes.
-The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al.
-2020).
-Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe 2011).
-Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Zhao, Yan, and Lu 2021; Zheng et al.
-2023; Chen et al.
-2025).
+The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies [mohsenvand2020contrastive].
+Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components [zander2011towards].
+Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data [zhao2021plug, zheng2023scatterformer, chen2025long].
 Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models.
@@
 These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability.
-For multiple tasks, two large EEG models, EEGPT (Wang et al.
-2024) and LaBraM (Jiang, Zhao, and Lu 2024) have recently been proposed.
-Both models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks.
-However, LaBraM uses a substantially larger dataset, and EEGPT’s performance across diverse specific tasks still leaves room for improvement.
-In this paper, we introduce the Diffusion Model (DM) (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM.
+For multiple tasks, three large EEG models, EEGPT [wangeegpt], LaBraM [jiang2024large], and CBraMod [wang2024cbramod] have recently been proposed.
+These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks.
+However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
+In this paper, we introduce the Diffusion Model (DM) [ho2020denoising, dhariwal2021diffusion] into EEG representation learning and propose a novel self-supervised framework called EEGDM.
 Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics.
-Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution.
+Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution [ho2020denoising].
 This makes DM potentially valuable for generic EEG representation learning.
@@
 This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
-We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al.
-2013; Cole and Voytek 2019).
+We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity [gramfort2013time, cole2019cycle].
 To address these challenges, we incorporate two key technical components in our EEGDM framework.
 First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information.
-Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers 2013).
+Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones [sanei2013eeg].
 This transformation mitigates the challenge of reconstructing the original noisy signal.
-Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al.
-2022) and Diffusion Transformers (DiTs) (Peebles and Xie 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space.
+Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) [rombach2022high] and Diffusion Transformers (DiTs) [peebles2023scalable]) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space.
 This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks.
-Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks with modest pre-training data requirements and exhibiting strong generalizability.
+Our experiments show that EEGDM effectively learns robust representations, achieving competitive performance across diverse downstream tasks and exhibiting strong generalizability.
 Our contributions are summarized as follows:
   • We propose EEGDM, a novel diffusion model-based self-supervised framework for EEG representation learning.
-By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics, even with limited pre-training data.
+By leveraging EEG signal generation as the self-supervised objective, EEGDM transforms the diffusion model into a powerful representation learner that captures rich EEG semantics.
   • EEGDM incorporates channel augmentation and elaborately designed PCA-based latent space operations to (1) enhance the conditional information available for DM, (2) enable robust EEG signal reconstruction, and (3) improve the cross-dataset generalizability of learned EEG representations.
-  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (a) competitive performance compared to existing methods and (b) strong cross-dataset transferability and practical utility in EEG applications.
-Related Work
-Self-supervised EEG Representation Learning
+  • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (1) competitive performance compared to existing methods and (2) strong cross-dataset transferability and practical utility in EEG applications.
+The remainder of this work is organized as follows.
+First, we describe the related works in Section II.
+Next, the proposed algorithm is introduced in detail in Section III.
+After that, experimental results are shown in Section IV.
+Finally, the conclusion is given in Section V.
+II Related work
+II-A Self-supervised EEG Representation Learning
 Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations.
-This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova 2019; He et al.
-2022).
-However, its full potential in EEG analysis remains largely unexplored (Jiang et al.
-2025).
-Some studies (Mohsenvand et al.
-2020; Banville et al.
-2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al.
-2020) to time series data.
-Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al.
-2024).
-For instance, BENDR (Kostas et al.
-2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part.
-LaBraM (Jiang, Zhao, and Lu 2024) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook.
-EEGPT (Wang et al.
-2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
+This methodology has led to significant breakthroughs in natural language processing and computer vision [kenton2019bert, he2022masked].
+However, its full potential in EEG analysis remains largely unexplored [jiang2024neurolm].
+Some studies [mohsenvand2020contrastive, banville2021uncovering] have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework [chen2020simple] to time series data.
+Another line of work extends the mask idea to EEG analysis [mohammadi2024eeg2rep].
+For instance, Chien et al.
+[chien2022maeeg] propose a masked auto-encoder, which learns the EEG representation by learning to reconstruct the masked EEG features using a Transformer architecture.
+BENDR [kostas2021bendr] uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part.
+LaBraM [jiang2024large] pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook.
+EEGPT [wangeegpt] introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
+CBraMod [wang2024cbramod] introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
 Unlike the above approaches, we explore self-supervised EEG representation learning via generative models.
 Our proposed EEGDM not only facilitates high-quality EEG generation but also enables the learning of rich and robust representations for downstream tasks.
-Diffusion Model
+II-B Diffusion Model
 Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics.
@@
 The pioneering work by Sohl-Dickstein et al.
-(Sohl-Dickstein et al.
-2015) and subsequent advances by Ho et al.
-(Ho, Jain, and Abbeel 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al.
-2022; Peebles and Xie 2023; Zhang, Rao, and Agrawala 2023).
-More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al.
-2023; Chen et al.
-2024b; Hudson et al.
-2024).
-In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al.
-2024; Chen et al.
-2024a; Huang, Li, and Chen 2025), or a few specific supervised tasks (Kim et al.
-2023, 2024).
+[sohl2015deep] and subsequent advances by Ho et al.
+[ho2020denoising] established diffusion models as powerful tools, leading to widespread applications in generative tasks [rombach2022high, peebles2023scalable, zhang2023adding].
+More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features [mittal2023diffusion, chen2024deconstructing, hudson2024soda].
+In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation [huang2024eegdfus, chen2024improving, huang2025sad], or a few specific supervised tasks [kim2023diff, kim2024brain].
 However, the potential of diffusion models for generic EEG representation learning remains largely unexplored.
@@
 In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
-Method
-Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying properties, structures, and dynamics of the data during the generative process.
-Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
 Figure 1: The overall architecture of EEGDM.
@@
 It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
-Framework Architecture
+III Method
+Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying properties, structures, and dynamics of the data during the generative process.
+Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
+III-A Framework Architecture
 Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation.
@@
 To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR.
-The overall architecture of EEGDM is illustrated in Figure 1.
+The overall architecture of EEGDM is illustrated in Fig.
+1.
 Specifically, given the original multi-channel EEG signals $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of EEG electrodes (channels) and $T$ is the total timestamp, we segment $X$ into ${\lfloor{\frac{T-t^{s}}{s^{t}}}\rfloor}+1$ samples.
@@
 $t^{s}$ per sample and a stride $s^{t}$.
-As shown in Figure 1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$.
+As shown in Fig.
+1, each multi-channel sample $x\in\mathbb{R}^{C\times{t^{s}}}$ is processed by PCA to obtain the latent representation $z$.
 Our diffusion model operates on this latent space, where noise is added and predicted.
@@
 Figure 2: The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
-The EEG Encoder
+III-B The EEG Encoder
 The representation output from the EEG encoder lies at the core of EEGDM, which (1) acts as the communication bridge between the encoder and the denoiser ($\epsilon_{\theta}$), and (2) guides the denoising process.
@@
 This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser.
-Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov 2006; Higgins et al.
-2017), our strategy enables the encoder to learn a representation that supports latent space denoising.
+Rather than reconstruction from scratch as most conventional denoisers do [hinton2006reducing, higgins2017beta], our strategy enables the encoder to learn a representation that supports latent space denoising.
 This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG.
 Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
-Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy 2021) that operates on the sequences of tokens.
+Our EEG encoder is based on the Vision Transformer (ViT) architecture [dosovitskiy2020image] that operates on the sequences of tokens.
 We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences.
@@
 EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data.
-Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al.
-2020) to be the most effective for extracting useful features.
+Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work [mohsenvand2020contrastive] to be the most effective for extracting useful features.
 Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$).
@@
 In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation.
-As shown in Figure 2, we train a linear classifier as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ cross-entropy loss.
-The implementation details can be found in the supplementary material.
-EEG Data Generation by Diffusion Model
+As shown in Fig.
+2, we train a linear predictor as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ task-specific loss functions for different downstream objectives.
+The implementation details can be found in the subsection IV-B.
+III-C EEG Data Generation by Diffusion Model
 As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging.
 To address this, we first enhance the SNR of the raw EEG signals using PCA.
-Similar to the LDMs (Rombach et al.
-2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals.
+Similar to the LDMs [rombach2022high], we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals.
 In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process.
 After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
-Diffusion Formulation
+III-C1 Diffusion Formulation
 The input to our proposed latent diffusion model, which is conditioned on channel augmentation, is the latent representation $z=Px$ obtained through PCA pre-processing.
@@
 To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al.
-(2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$.
+[nichol2021improved]: train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$.
 After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction.
@@
 $\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$. (4)
-Conditional Guidance
+III-C2 Conditional Guidance
 We incorporate the encoded EEG representation as conditional input for the diffusion model.
@@
 When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$.
-Under this setting, classifier-free guidance (Ho and Salimans 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized.
+Under this setting, classifier-free guidance [ho2022classifier] can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized.
 According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$.
@@
 The classifier-free guidance sampling method can significantly improve the quality of generated samples by effectively utilizing the conditional information we introduced.
-Methods Model Size Balanced Accuracy ↑ AUROC ↑
+III-C3 Conditional Input Setup
+The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models.
+We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$).
+Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$.
+Additionally, we regress the dimension-wise scaling parameters $\alpha$, which are applied just before each residual connection in the DiT blocks.
+The MLP for all $\alpha$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function.
+In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
+IV Experiment
+IV-A Datasets
+Pre-training datasets.
+We utilize the Temple University Hospital EEG Corpus (TUEG) [obeid2016temple] to pre-train the proposed EEGDM.
+TUEG is a very large publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments.
+It has over 40 different channel configurations and varying temporal durations of recordings.
+The majority of the recordings are acquired at a sampling rate of 256 Hz.
+TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
+Datasets of downstream specific tasks.
+(1) TUEV (event type classification) [harati2015improved]: An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
+The EEG signals are recorded at 23 channels and 256 Hz sampling rate.
+(2) TUAB (abnormal detection) [lopez2015automated]: A corpus of EEG recordings annotated as either normal or abnormal.
+Similar to TUEV, the EEG signals of TUAB are recorded at 23 channels and 256 Hz sampling rate.
+(3) SHU-MI (motor imagery classification) [ma2022large]: An EEG motor imagery dataset designed for decoding imagined movements.
+It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand.
+(4) CL-Drive (cognitive load assessment) [angkan2024multimodal]: It is a multimodal dataset designed for cognitive load assessment, which includes EEG signals.
+The EEG data are recorded using four channels at a sampling frequency of 256 Hz.
+This study explores the ternary classification of cognitive load into high, medium, and low levels.
+(5) Seed-VII (emotion recognition) [jiang2024seed]: It is a large-scale multimodal emotion dataset.
+Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise).
+The preprocessed EEG signals were downsampled to 200 Hz.
+(6) DMER (emotion distribution prediction) [yang2024multimodal]: It is a multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos).
+In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) [mackinnon1999short], reporting the intensity of ten emotional adjectives.
+Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution.
+The EEG recordings comprise 18 channels sampled at 300 Hz.
+It is worth noting that we primarily select downstream datasets from the past five years, a strategic shift from the benchmarks used in works like CBraMod [wang2024cbramod] or EEGPT [wangeegpt].
+This decision stems from the need to evaluate the model against the fidelity and complexity of modern neuroscience standards.
+Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms — ranging from naturalistic stimuli to multimodal retrieval — better assess the capacity for generalizable, high-level cognitive decoding.
+However, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets.
+Despite falling outside the five-year window, these two widely adopted benchmarks are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
+IV-B Experimental Setup
+For the pre-training dataset TUEG, following prior work [wang2024cbramod], we select 19 common EEG channels.
+All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels × 6,000 time points).
+Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample.
+For all downstream datasets, we maintain consistency with our pre-training data by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
+For the downstream datasets TUEV and TUAB, consistent with previous studies [wangeegpt, jiang2024large, wang2024cbramod], we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively.
+We adopt the train-test splits provided by these datasets.
+For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold subject-wise cross-validation.
+For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples.
+We employ a subject-independent split strategy for these datasets.
+We fine-tune EEGPT [wangeegpt], LaBraM-Base [jiang2024large] (the only variant with available pre-trained weights), and CBraMod [wang2024cbramod] based on their publicly available code and pre-trained weights, except when their experimental results are already reported in the original papers.
+For the EEG generator (diffusion module) and encoder, we adopt the DiT and ViT architectures, respectively.
+Following common practices in generative modeling literature, we maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999 during pre-training.
+The specific hyperparameters for pre-training are detailed in Table I.
+We evaluate the representation learning capabilities of EEGDM through fine-tuning on downstream tasks.
+Specifically, we predict corresponding labels using a linear predictor based on the outputs of the EEG encoder.
+Following the common evaluation protocols, we present results under the fine-tuning setting, where the predictor is jointly trained with the EEG encoder (i.e., the pre-trained encoder is not frozen).
+We employ cross-entropy loss for downstream classification tasks and KL divergence loss for downstream distribution prediction tasks.
+The hyperparameters for fine-tuning EEGDM on downstream tasks are provided in Table II.
+We conduct a series of quantitative and qualitative experiments to evaluate the proposed model, demonstrating its strong representation capabilities across diverse downstream datasets.
+TABLE I: Hyperparameters for EEGDM pre-training.
+Hyperparameter Value
+Optimization Batch size 32
+Peak learning rate 5e-4
+Minimal learning rate 5e-5
+Learning rate scheduler OneCycleLR
+Weight decay 0.05
+EMA decay rate 0.9999
+Gradient clipping norm 1
+Optimizer AdamW
+${\beta}_{1}$ 0.9
+${\beta}_{2}$ 0.999
+EEG Generator Layers N 8
+Hidden size 512
+Heads 8
+Normalization type LayerNorm
+EEG Encoder Layers 8
+Hidden size 512
+MLP size 2048
+Heads 8
+Sampling Classifier-free guidance 9
+Diffusion training steps 1000
+TABLE II: Hyperparameters for downstream fine-tuning. ($\star$) Depends on the dataset.
+Hyperparameter Value
+Batch size 64
+Peak learning rate 5e-4
+Minimal learning rate 5e-5
+Learning rate scheduler OneCycleLR
+Optimizer AdamW
+${\beta}_{1}$ 0.9
+${\beta}_{2}$ 0.999
+Weight decay 0.05
+Total epochs 30, 150
+Warmup Epochs 6, 30
+IV-C Results and Analysis
+In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis.
+All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics.
+The reported model size corresponds to the downstream networks used for fine-tuning, as the generative module is not involved in downstream tasks.
+Results are compared against baselines, as shown in Tables III, IV, V and VI.
+Compared to other approaches, EEGDM consistently maintains either the best or second-best performance in key metrics (balanced accuracy or KL divergence).
+TABLE III: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
+Methods Params Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
 […]
@@
 […]
-Table 1: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
-Methods Model Size Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
+[…]
+[…]
+TABLE IV: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
+Methods Params Balanced Accuracy ↑ AUROC ↑ AUC-PR ↑
 […]
@@
 […]
-Table 2: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
-Datasets Methods Balanced Accuracy ↑ Cohen’s Kappa ↑ Weighted F1 / AUROC ↑
+[…]
+[…]
+TABLE V: Performance comparison on DMER. The best metric is highlighted in bold, and the second-best metric is underlined.
+Methods Chebyshev ↓ Clark ↓ Canberra ↓ KL ↓ Cosine ↑ Intersection ↑
+[…]
+[…]
+[…]
+[…]
+TABLE VI: The results of different methods on various datasets. The best metric is highlighted in bold, and the second-best metric is underlined.
+Datasets Methods Balanced Accuracy ↑ Weighted F1 / AUROC ↑ Cohen’s Kappa / AUC-PR ↑
 […]
@@
 […]
-Table 3: Performance comparison on BCIC. The best metric is highlighted in bold, and the second-best metric is underlined.
-Variants PCA EEG Augmentation BCIC-2A ↑ BCIC-2B ↑ TUAB (10%) ↑ TUEV (10%) ↑
 […]
@@
 […]
-Table 4: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
-Conditional Input Setup
-The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models.
-We adopt the adaLN-Zero block from DiT, which has shown superior performance, to handle the conditional inputs (the EEG representations $e$ and the noise timestep $t$).
-Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$.
-Additionally, we regress the dimension-wise scaling parameters $\alpha$, which are applied just before each residual connection in the DiT blocks.
-The MLP for all $\alpha$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function.
-In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
-Experiments
-Datasets
-Pre-training datasets.
-We pre-train the proposed EEGDM on several publicly available EEG datasets covering a range of paradigms, including PhysioMI (motor imagery and execution dataset) (Goldberger et al.
-2000), TSU (steady-state visual evoked potential dataset) (Wang et al.
-2016), SEED (emotion recognition dataset) (Zheng and Lu 2015), and M3CV (multi-subject multi-session multi-paradigm dataset) (Huang et al.
-2022).
-Datasets of downstream specific tasks.
-(1) TUAB (abnormal detection) (Obeid and Picone 2016): A corpus of EEG recordings annotated as either normal or abnormal.
-(2) TUEV (event type classification) (Obeid and Picone 2016): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
-(3) BCIC-2A (motor imagery classification) (Tangermann et al.
-2012): A four-class motor imagery dataset comprising left hand, right hand, feet, and tongue movement imagination tasks.
-(4) BCIC-2B (motor imagery classification) (Steyrl et al.
-2016): A binary motor imagery dataset containing left-hand and right-hand movement imagination tasks.
-Experimental Setup
-We evaluate EEGDM’s representation learning capabilities through fine-tuning on downstream tasks.
-Following previous work (Wang et al.
-2024), we adopt standard preprocessing procedures including resampling, normalization, segmentation, and flattening.
-For TUAB and TUEV datasets, we use the provided train-test splits, while BCIC-2A and BCIC-2B employ Leave-One-Subject-Out (LOSO) cross-validation.
-The reported model sizes correspond to the downstream networks used for fine-tuning, as the generative modules are not involved in downstream tasks.
-We compare with state-of-the-art baselines and provide results for both EEGDM (25M parameters) and a compact variant, EEGDM-Tiny (3M parameters), which uses a smaller pre-training architecture.
-Detailed configurations for both model variants are provided in the supplementary material.
-Complete implementation details and hyperparameters for both pre-training and fine-tuning are also provided in the supplementary material.
-Results and Analysis
-In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis.
-All experiments are conducted with three random seeds, and we report mean and standard deviation values as shown in Tables 1, 2, and 3.
-Additional results under different fine-tuning settings are provided in the supplementary material.
-On the TUAB dataset, EEGDM achieves a balanced accuracy of 80.17%, outperforming both the previous state-of-the-art method EEGPT (79.83%) and other baselines.
-The AUROC score of 87.39% further demonstrates strong discriminative capability.
-For TUEV classification, EEGDM obtains 60.81% balanced accuracy with a Cohen’s Kappa coefficient 0.5218, achieving either (1) good balance between performance and pre-training data size, when compared to EEGPT, or (2) substantial improvements over other baselines.
-On motor imagery classification, EEGDM shows competitive performance, achieving 55.95% balanced accuracy on BCIC-2A and 72.19% on BCIC-2B, with the latter outperforming both EEGPT and LaBraM.
-Regarding the comparison with EEGPT and LaBraM, the following detailed explanations are necessary.
-Following EEGPT (Wang et al.
-2024), all comparisons are conducted with LaBraM-Base (Jiang, Zhao, and Lu 2024).
-The LaBraM results are based on the same pre-training dataset as ours, rather than their original larger dataset (which includes TUEG dataset, whose distribution is similar to the TUAB and TUEV).
-Remarkably, our EEGDM achieves performance comparable to or better than EEGPT despite using a smaller pre-training dataset.
-Even our compact EEGDM-Tiny model maintains competitive performance while using substantially fewer parameters.
-These results validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
-Ablation Study
-We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table 4, in which the average balanced accuracy on three random seeds is reported.
-For BCIC-2A and BCIC-2B, we report the results using the LOSO evaluation protocol.
-Variant D consistently achieves the best performance across all settings, demonstrating that both PCA and EEG augmentation are important for our model.
-Furthermore, for TUAB and TUEV, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last two columns in Table 4); we also use this setting in Figure 3.
+[…]
+[…]
+TABLE VII: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
+Variants PCA EEG Augmentation TUEV TUAB (10%)
+[…]
+[…]
+[…]
+[…]
+On the TUEV dataset, EEGDM achieves a balanced accuracy of 68.19%, outperforming both the previous state-of-the-art method CBraMod (66.71%) and other baselines.
+This superiority highlights the exceptional discriminative power of the representations learned by the diffusion model for classifying different event types in EEG signals.
+Table IV shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset.
+Securing the second-best performance on those metrics, EEGDM overall demonstrates comparable performance with a competitive margin.
+The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well for binary abnormal detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
+Table V presents the results for the DMER dataset, which uses distributional metrics.
+EEGDM achieves best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second.
+The overall high performance on DMER, a task requiring the prediction of a continuous emotion distribution rather than a discrete class, strongly suggests that the diffusion-based pre-training learns rich and fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in the EEG signals.
+Table VI further validates the generalizability and robustness of EEGDM across diverse BCI tasks.
+Specifically, on the SHU-MI dataset, EEGDM ranks second in terms of balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod.
+For the CL-Drive dataset, EEGDM shows the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities.
+On the Seed-VII dataset, EEGDM continues its strong performance with the best performance in balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
+These comprehensive results across diverse datasets validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
+(a) TUEV
+(b) TUAB
+Figure 3: Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
+IV-D Ablation Study
+We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table VII, in which the average balanced accuracy on three random seeds is reported.
+Variant D consistently achieves the best performance across all settings, demonstrating that PCA and EEG augmentation are important for our model.
+However, we find that using PCA or EEG augmentation alone is either detrimental to or has minimal impact on our model’s performance.
+This may be attributed to the fact that PCA reduces the dimensionality of the source signal, making the model more sensitive to the diversity of conditional information, while data augmentation provides sufficient diversity to support learning in this reduced-dimensional space.
+Consequently, the PCA and EEG augmentation operations achieve complementary effects in terms of generalization performance and accuracy.
+Furthermore, for TUAB, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last column in Table VII); we also use this setting in Fig.
+3.
 The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
 We also investigate the impact of the number of PCA components on downstream task performance.
-Figure 3 presents results on TUEV and BCIC-2B with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance.
+Fig.
+3 presents results on TUEV and TUAB with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance.
 In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance.
 To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
-To evaluate generation quality, we use the diffusion model in EEGDM under the same pre-training configuration, and compute Pearson correlation coefficients between the generated and original signals in both time and frequency domains.
-When PCA is applied, the correlation coefficients improve significantly to 0.36 (time domain) and 0.88 (frequency domain), compared to lower values of 0.003 and 0.83 without PCA.
-These results demonstrate that integrating PCA substantially improves the quality of EEG data generation, which further facilitates better representation learning.
-Figure 3: Performance comparison with different numbers of PCA components on TUEV and BCIC-2B datasets.
-Visualization
-To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton 2008) as presented in Figure 4.
-We randomly sample EEG signals from the test sets of four downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder.
+(a) TUEV
+(b) TUAB
+Figure 4: The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
+Figure 5: Visualization of reconstructed EEG signals and their PSD.
+IV-E Visualization
+To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE [van2008visualizing] as presented in Fig.
+4.
+We randomly sample EEG signals from the test sets of two downstream tasks in a class-balanced manner and extract their representations using our pre-trained encoder.
 The t-SNE visualizations exhibit well-separated clusters across all datasets, indicating that our EEGDM learns discriminative representations capable of capturing task-relevant patterns.
 This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks.
-Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses provided in the supplementary material.
-(a) TUAB
-(b) TUEV
-(c) BCIC-2A
-(d) BCIC-2B
-Figure 4: The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
-In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by the visualizations in supplementary material; (2) Effective representation learning, evidenced by distinct class separation in t-SNE visualizations and consistent performance improvements across tasks; (3) Competitive performance across diverse downstream tasks using modest pre-training data sizes, particularly given that EEGDM matches or exceeds EEGPT’s performance while using smaller pre-training datasets.
-Conclusion
+Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses.
+As shown in Fig.
+5, the reconstructed EEG signals on the pre-training dataset, as well as their frequency-domain representations, closely match the ground truth.
+While minor discrepancies remain at the fine-grained level, the global temporal structure and frequency characteristics are well preserved.
+These results indicate that the proposed model can generate realistic EEG signals and effectively capture high-level semantic information from EEG data.
+In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by signal reconstruction visualization; (2) Effective representation learning, as demonstrated by distinct class separation in t-SNE visualization; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent performance improvements across tasks.
+V Conclusion
 We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation.

### v3

@@
 EEGDM: Learning EEG Representation with Latent Diffusion Model
-Abstract
+Abstract.
 Recent advances in self-supervised learning for EEG representation have largely relied on masked reconstruction, where models are trained to recover randomly masked signal segments.
@@
 Experimental results show that EEGDM (1) reconstructs high-quality EEG signals, (2) learns robust representations, and (3) achieves competitive performance across diverse downstream tasks, thus exploring a new direction for self-supervised EEG representation learning.
-Index Terms:
-Self-supervised learning, diffusion model, EEG representation learning.
-I Introduction
+Electroencephalography, Representation Learning, Self-supervised Learning, Diffusion Transformers
+conference: ; ; copyright: noneisbn: 978-1-4503-XXXX-X/2018/06submissionid: 5469ccs: Computing methodologies Machine learningccs: Human-centered computing Human computer interaction (HCI)$c$s: Information systems Multimedia information systems
+1. Introduction
 Electroencephalography (EEG) is a non-invasive technique that measures spontaneous electrical activity in the brain through scalp electrodes.
-The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies [mohsenvand2020contrastive].
-Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components [zander2011towards].
-Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data [zhao2021plug, zheng2023scatterformer, chen2025long].
+The interpretation and utilization of EEG signals generally require sophisticated signal processing pipelines and advanced feature extraction methodologies (Mohsenvand et al., 2020).
+Conventional EEG signal analysis typically employs manually engineered operations tailored to specific EEG features of interest (e.g., band-pass filtering in a specific frequency range), potentially excluding other informative signal components (Zander and Kothe, 2011).
+Recently, deep learning has introduced a new research paradigm for EEG signal analysis due to its capability to automatically learn efficient representations from original data (Jia et al., 2020, 2021; Jiang et al., 2024a).
 Current deep learning-based approaches, which predominantly rely on large-scale EEG datasets, can be broadly classified into two categories: (1) task-specific models, and (2) multi-task large models.
@@
 These discrepancies necessitate model training on individual datasets, thereby severely limiting cross-dataset generalizability.
-For multiple tasks, three large EEG models, EEGPT [wangeegpt], LaBraM [jiang2024large], and CBraMod [wang2024cbramod] have recently been proposed.
+For multiple tasks, three large EEG models, EEGPT (Wang et al., 2024), LaBraM (Jiang et al., 2024c), and CBraMod (Wang et al., 2025) have recently been proposed.
 These models learn generic representations through masked reconstruction pre-training objective before being fine-tuned for various downstream tasks.
 However, masked reconstruction has limited capacity in capturing the global dynamics and long-term dependencies essential for characterizing neural activity.
-In this paper, we introduce the Diffusion Model (DM) [ho2020denoising, dhariwal2021diffusion] into EEG representation learning and propose a novel self-supervised framework called EEGDM.
+In this paper, we introduce the Diffusion Model (DM) (Ho et al., 2020; Dhariwal and Nichol, 2021) into EEG representation learning and propose a novel self-supervised framework called EEGDM.
 Our framework leverages EEG signal synthesis as a self-supervised objective, turning the diffusion model into a strong representation learner capable of capturing EEG semantics.
-Unlike popular masked reconstruction approaches that primarily focus on local feature recovery, DM learns effective representations through a self-supervised denoising objective that requires understanding the underlying data distribution [ho2020denoising].
+Unlike popular masked reconstruction paradigms that primarily focus on local patches recovery, DM learns effective representations through a self-supervised denoising objective.
+This objective necessitates a profound understanding of the underlying data distribution, thereby compelling the latent space to preserve macro-level neural patterns.
+To successfully denoise a latent representation into a coherent EEG segment, the model must capture the holistic phase relationships and cross-channel correlations that define specific brain states.
 This makes DM potentially valuable for generic EEG representation learning.
@@
 This design enables the encoder to learn the effective and generic representations inherent in EEG signals that are crucial for downstream tasks.
-We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity [gramfort2013time, cole2019cycle].
-To address these challenges, we incorporate two key technical components in our EEGDM framework.
+We note that EEG signals present unique challenges due to their inherently low signal-to-noise ratio (SNR) and complex characteristics, including stochasticity, nonstationarity, and nonlinearity (Gramfort et al., 2013; Cole and Voytek, 2019).
+To address these challenges, we incorporate two synergistic components into our EEGDM framework.
 First, we employ channel augmentation that provides diverse training perspectives while enriching the conditional information.
-Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones [sanei2013eeg].
+Second, we utilize Principal Component Analysis (PCA) to project EEG signals into a latent space with an improved SNR, as it enhances signal-relevant components while suppressing noise-related ones (Sanei and Chambers, 2013).
 This transformation mitigates the challenge of reconstructing the original noisy signal.
-Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) [rombach2022high] and Diffusion Transformers (DiTs) [peebles2023scalable]) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space.
+More importantly, PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views.
+Meanwhile, the diverse augmented inputs prevent the model from losing critical information when training against a reduced-dimension (PCA) target.
+Inspired by the success of state-of-the-art DMs (e.g., Latent Diffusion Models (LDMs) (Rombach et al., 2022) and Diffusion Transformers (DiTs) (Peebles and Xie, 2023)) operating in latent spaces, we apply our DM to this PCA-derived latent space rather than the original noisy EEG space.
 This enables EEGDM to focus on learning meaningful signal patterns for effective representation learning across diverse downstream tasks.
@@
   • We validate the effectiveness and generalizability of EEGDM across multiple EEG analysis tasks, demonstrating its (1) competitive performance compared to existing methods and (2) strong cross-dataset transferability and practical utility in EEG applications.
-The remainder of this work is organized as follows.
-First, we describe the related works in Section II.
-Next, the proposed algorithm is introduced in detail in Section III.
-After that, experimental results are shown in Section IV.
-Finally, the conclusion is given in Section V.
-II Related work
-II-A Self-supervised EEG Representation Learning
+2. Related Work
+2.1. Self-supervised EEG Representation Learning
 Self-supervised learning can leverage large amounts of unlabeled data for pre-training, thereby reducing the reliance on manual annotations.
-This methodology has led to significant breakthroughs in natural language processing and computer vision [kenton2019bert, he2022masked].
-However, its full potential in EEG analysis remains largely unexplored [jiang2024neurolm].
-Some studies [mohsenvand2020contrastive, banville2021uncovering] have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework [chen2020simple] to time series data.
-Another line of work extends the mask idea to EEG analysis [mohammadi2024eeg2rep].
-For instance, Chien et al.
-[chien2022maeeg] propose a masked auto-encoder, which learns the EEG representation by learning to reconstruct the masked EEG features using a Transformer architecture.
-BENDR [kostas2021bendr] uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part.
-LaBraM [jiang2024large] pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook.
-EEGPT [wangeegpt] introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
-CBraMod [wang2024cbramod] introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
+This methodology has led to significant breakthroughs in natural language processing and computer vision (Kenton and Toutanova, 2019; He et al., 2022).
+However, its full potential in EEG analysis remains largely unexplored (Jiang et al., 2025).
+Some studies (Mohsenvand et al., 2020; Banville et al., 2021) have explored learning representations of EEG signals via contrastive learning, which trains a channel feature extractor by extending the SimCLR framework (Chen et al., 2020) to time series data.
+Another line of work extends the mask idea to EEG analysis (Mohammadi Foumani et al., 2024).
+For instance, BENDR (Kostas et al., 2021) uses a convolutional encoder to extract EEG features from local time windows, masks a portion of them, and utilizes a Transformer to predict the information in the masked part.
+LaBraM (Jiang et al., 2024c) pre-trains a Transformer model by predicting the original neural codes of masked EEG channel patches based on a defined neural codebook.
+EEGPT (Wang et al., 2024) introduces the alignment of spatio-temporal representation, implemented by a mask-based dual self-supervised learning for efficient extraction of EEG features.
+CBraMod (Wang et al., 2025) introduces a structurally aware EEG foundation model that separately captures heterogeneous spatial–temporal dependencies through a criss-cross transformer and achieves strong cross-dataset generalization via large-scale masked pre-training.
 Unlike the above approaches, we explore self-supervised EEG representation learning via generative models.
 Our proposed EEGDM not only facilitates high-quality EEG generation but also enables the learning of rich and robust representations for downstream tasks.
-II-B Diffusion Model
+2.2. Diffusion Model
 Diffusion models learn complex data distributions through a progressive denoising process that requires a deep understanding of the underlying data structure and semantics.
@@
 The pioneering work by Sohl-Dickstein et al.
-[sohl2015deep] and subsequent advances by Ho et al.
-[ho2020denoising] established diffusion models as powerful tools, leading to widespread applications in generative tasks [rombach2022high, peebles2023scalable, zhang2023adding].
-More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features [mittal2023diffusion, chen2024deconstructing, hudson2024soda].
-In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation [huang2024eegdfus, chen2024improving, huang2025sad], or a few specific supervised tasks [kim2023diff, kim2024brain].
-However, the potential of diffusion models for generic EEG representation learning remains largely unexplored.
-This presents a significant opportunity, as learning the complex temporal and spatial characteristics of EEG signals can benefit from the understanding of rich data distributions inherent in diffusion models’ denoising process.
-In this paper, we present a novel method that exploits the representational capacity of diffusion models for self-supervised EEG representation learning.
-Figure 1: The overall architecture of EEGDM.
+(Sohl-Dickstein et al., 2015) and subsequent advances by Ho et al.
+(Ho et al., 2020) established diffusion models as powerful tools, leading to widespread applications in generative tasks (Rombach et al., 2022; Peebles and Xie, 2023; Zhang et al., 2023).
+More recently, their potential for representation learning has gained attention, with emerging studies showing promising results in learning discriminative features (Mittal et al., 2023; Chen et al., 2024b; Hudson et al., 2024).
+Figure 1.
+The overall architecture of EEGDM.
 The original EEG data is projected onto a latent space using PCA.
@@
 It is combined with the noise timestep representation and then used as conditional information to modulate the DiT block, thus guiding the denoising of the latent representation.
-III Method
-Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying properties, structures, and dynamics of the data during the generative process.
+In the EEG domain, existing diffusion-based approaches have primarily focused on generative applications, such as EEG synthesis and data augmentation (Huang et al., 2024; Chen et al., 2024a; Huang et al., 2025), or on a few specific supervised tasks (Kim et al., 2023, 2024).
+Recently, Puah et al.
+(Puah et al., 2025) proposed an EEG representation learning framework, also termed EEGDM.
+To avoid ambiguity, we refer to their method as Puah-EEGDM-SSM.
+Their approach integrates a Structured State-Space Model (SSM) into the diffusion pre-training paradigm to better capture the temporal dynamics of EEG signals.
+In contrast, our EEGDM adopts a fundamentally different architectural paradigm.
+Specifically, we employ a Latent Diffusion Transformer (DiT) operating on a refined latent manifold, which is specifically designed to capture global inter-channel spatial synchrony and complex phase relationships—critical neural features that are overlooked by the linear recurrence mechanisms inherent in SSM.
+3. Method
+Our study is motivated by the following observation: Diffusion models capable of generating rich and high-fidelity content from scratch are likely to effectively characterize the underlying distributions, structures, and dynamics of the data during the generative process.
 Therefore, we explore self-supervised EEG representation learning via the diffusion model, and design the following framework architecture.
-III-A Framework Architecture
-Our proposed self-supervised EEG representation learning framework, EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation.
+3.1. Framework Architecture
+Our proposed self-supervised EEG representation learning framework, named EEGDM, is formulated as a latent diffusion model (DM) conditioned on EEG channel augmentation.
 Our goal is to generate high-quality EEG signals while equipping the model with a generalizable understanding of EEG data that can be effectively adapted to various downstream tasks.
-It is worth noting that in the denoising process of DM, directly reconstructing raw EEG signals is challenging due to their inherently low SNR.
-To address this, we utilize PCA to preprocess the original EEG for noise reduction, and perform the reconstruction on the preprocessed EEG data with a relatively high SNR.
-The overall architecture of EEGDM is illustrated in Fig.
+Notably, reconstructing raw EEG signals directly remains formidable due to their inherently low SNR.
+To mitigate this, we employ PCA to project raw data into a refined latent space with enhanced SNR, shifting the reconstruction target to these denoised components.
+By focusing on the top-K principal components, the diffusion module effectively constrains the search space to the most energetically significant neural manifolds, thereby alleviating the optimization burden.
+Crucially, while EEG augmentation introduces view diversity to force the extraction of invariant features, PCA provides a stable, denoised reconstruction objective.
+The overall architecture of EEGDM is shown in Fig.
 1.
@@
 Consequently, this enhances the model’s representation learning capability.
-Figure 2: The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
-III-B The EEG Encoder
+3.2. The EEG Encoder
 The representation output from the EEG encoder lies at the core of EEGDM, which (1) acts as the communication bridge between the encoder and the denoiser ($\epsilon_{\theta}$), and (2) guides the denoising process.
@@
 This strategy encourages the encoder to capture high-level semantic information in the EEG signals, while leaving the reconstruction of local details to the denoiser.
-Rather than reconstruction from scratch as most conventional denoisers do [hinton2006reducing, higgins2017beta], our strategy enables the encoder to learn a representation that supports latent space denoising.
+Rather than reconstruction from scratch as most conventional denoisers do (Hinton and Salakhutdinov, 2006; Higgins et al., 2017), our strategy enables the encoder to learn a representation that supports latent space denoising.
 This alleviates the burden of compressing all signal information into the representation, allowing it to prioritize the most distinctive and informative features of the EEG.
 Such a representation is particularly valuable for adapting to a wide range of downstream EEG tasks.
-Our EEG encoder is based on the Vision Transformer (ViT) architecture [dosovitskiy2020image] that operates on the sequences of tokens.
+Our EEG encoder is based on the Vision Transformer (ViT) architecture (Dosovitskiy, 2021) that operates on the sequences of tokens.
 We segment each EEG channel into tokens and then flatten the tokens to form consistent biosignal sequences.
@@
 EEG augmentations are a set of transformations that preserve the semantic information in the EEG channels without altering the interpretation of the EEG data.
-Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work [mohsenvand2020contrastive] to be the most effective for extracting useful features.
+Here, we employ two augmentations (zero-masking and amplitude scaling) which were shown in prior work (Mohsenvand et al., 2020) to be the most effective for extracting useful features.
 Specifically, we use a $\omega$-length window without overlap to segment each EEG channel sample into patches ($x^{\prime}=\{x^{\prime}_{j,k}\in\mathbb{R}^{\omega}|j=1,2,\cdots,C,k=1,2,\cdots,{\lfloor{\frac{t^{s}}{\omega}}\rfloor}\}$).
@@
 We use average pooling on the output embeddings to obtain the EEG representation, which is then combined with the noise timestep representation to guide the denoising of the diffusion model.
-In addition, we perform downstream EEG tasks by using task-specific prediction heads on the EEG representation.
-As shown in Fig.
-2, we train a linear predictor as the prediction head based on the pre-trained encoder in fine-tuning and then analyze the model’s performance on the downstream tasks, in which we employ task-specific loss functions for different downstream objectives.
-The implementation details can be found in the subsection IV-B.
-III-C EEG Data Generation by Diffusion Model
+As illustrated in Fig.
+2, we evaluate the model’s performance on various downstream tasks by appending a task-specific linear prediction head to the pre-trained encoder.
+For each downstream objective, tailored loss functions are employed to ensure optimal task-specific adaptation.
+The implementation details are provided in Section 4.2.
+Figure 2. The pre-trained encoder is connected to a linear prediction head to perform downstream tasks.
+3.3. EEG Data Generation by Diffusion Model
 As aforementioned, due to the inherently low SNR of original EEG signals, accurate reconstruction of them directly in the denoising process is challenging.
-To address this, we first enhance the SNR of the raw EEG signals using PCA.
-Similar to the LDMs [rombach2022high], we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals.
+To address this, we utilize PCA to project raw data into a high-SNR subspace, establishing a refined latent manifold as the denoising objective.
+Similar to the LDMs (Rombach et al., 2022), we find a perceptually equivalent yet computationally more suitable space in which the diffusion model is trained for generating high-SNR EEG signals.
 In addition, we use the high-quality representation from the EEG encoder to guide the diffusion process.
 After denoising in the latent space, we apply inverse PCA to project the latent EEG representations back to the original signal space, thereby reconstructing the final EEG signals.
-III-C1 Diffusion Formulation
+3.3.1. Diffusion Formulation
 The input to our proposed latent diffusion model, which is conditioned on channel augmentation, is the latent representation $z=Px$ obtained through PCA pre-processing.
@@
 By applying the reparameterization trick, we can do the sampling as follows:
-$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$ (1)
+(1) $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},$
 where $\epsilon_{t}\sim{\mathcal{N}(0,\mathbf{I})}$. The diffusion model is trained to learn the reverse denoising process:
-$p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$ (2)
+(2) $p_{\theta}(z_{t-1}|z_{t})=\mathcal{N}(\mu_{\theta}(z_{t}),\Sigma_{\theta}(z_{t})),$
 where the network is used to predict the statistics of $p_{\theta}$.
 By reparameterizing $\mu_{\theta}$ as a noise prediction network (i.e., denoiser) $\epsilon_{\theta}$, the model can be trained using the mean squared error between the predicted noise $\epsilon_{\theta}(z_{t})$ and the ground truth sampled Gaussian noise $\epsilon_{t}$:
-$\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$ (3)
+(3) $\mathcal{L}_{simple}(\theta)={\left\|{\epsilon_{\theta}(z_{t})-\epsilon_{t}}\right\|}_{2}^{2}.$
 To train the diffusion model with a learned reverse process covariance $\Sigma_{\theta}$, we follow the approach of Nichol et al.
-[nichol2021improved]: train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$.
+(Nichol and Dhariwal, 2021): train $\epsilon_{\theta}$ with $\mathcal{L}_{simple}$ and train $\Sigma_{\theta}$ with $\mathcal{L}_{vlb}$.
 After the final DiT block, we use a standard linear decoder to decode the token sequence into an output noise prediction and an output diagonal covariance prediction.
@@
 Finally, we obtain the denoised latent representation $\widetilde{z}$, which is then projected back to the original signal space using inverse PCA to reconstruct the original EEG signals:
-$\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$. (4)
-III-C2 Conditional Guidance
+(4) $\widetilde{x}=P^{-1}\widetilde{z}=P^{T}\widetilde{z}$.
+3.3.2. Conditional Guidance
 We incorporate the encoded EEG representation as conditional input for the diffusion model.
@@
 When the conditional input is $e$, the reverse process of the diffusion model is $p_{\theta}(z_{t-1}|z_{t},e)$, where both $\epsilon_{\theta}$ and $\Sigma_{\theta}$ are conditioned on $e$.
-Under this setting, classifier-free guidance [ho2022classifier] can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized.
+Under this setting, classifier-free guidance (Ho and Salimans, 2022) can be applied to encourage the sampling process to find $z$ such that $\log p(e|z)$ is maximized.
 According to Bayes’ rule, $p(e|z)\propto{p(z|e)/p(z)}$, and thus $\nabla_{z}\log p(e|z)\propto{\nabla_{z}\log p(z|e)-\nabla_{z}\log p(z)}$.
 By interpreting the output of the diffusion model as a score function, the sampling process can be guided by:
-$\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$ (5)
+(5) $\widetilde{\epsilon}_{\theta}(z_{t},e)\propto{\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\epsilon_{\theta}(z_{t},e)-\epsilon_{\theta}(z_{t},\emptyset))}},$
 $\widetilde{\epsilon}_{\theta}(z_{t},e)=\epsilon_{\theta}(z_{t},\emptyset)+s\cdot{(\nabla_{z}\log p(z|e))},$
@@
 The classifier-free guidance sampling method can significantly improve the quality of generated samples by effectively utilizing the conditional information we introduced.
-III-C3 Conditional Input Setup
+3.3.3. Conditional Input Setup
 The treatment of conditional inputs and the initialization process significantly influence the generation quality in diffusion models.
@@
 Specifically, we regress the scale and shift parameters $\gamma$ and $\beta$ in the Transformer blocks based on the sum of the $e$ and $t$.
-Additionally, we regress the dimension-wise scaling parameters $\alpha$, which are applied just before each residual connection in the DiT blocks.
-The MLP for all $\alpha$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function.
+Additionally, we regress the dimension-wise scaling parameters $\rho$, which are applied just before each residual connection in the DiT blocks.
+The MLP for all $\rho$ is initialized to output the zero vector, which effectively initializes the entire DiT block as an identity function.
 In addition to layer normalization, we incorporate conditional embeddings into residual connections, which enhance the vanilla adaLN and improve the quality of the generated EEG signals.
-IV Experiment
-IV-A Datasets
+4. Experiment
+4.1. Datasets
 Pre-training datasets.
-We utilize the Temple University Hospital EEG Corpus (TUEG) [obeid2016temple] to pre-train the proposed EEGDM.
-TUEG is a very large publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments.
-It has over 40 different channel configurations and varying temporal durations of recordings.
-The majority of the recordings are acquired at a sampling rate of 256 Hz.
+We utilize the Temple University Hospital EEG Corpus (TUEG) (Obeid and Picone, 2016) to pre-train the proposed EEGDM.
+TUEG is a large-scale publicly available clinical EEG dataset containing clinical-grade recordings collected in real-world hospital environments.
+It encompasses over 40 different channel configurations and recordings of varying temporal durations, with the majority acquired at a sampling rate of 256 Hz.
 TUEG exhibits substantial heterogeneity in patient demographics, channel configurations, and recording equipment, making it particularly suitable for pre-training our model.
-Datasets of downstream specific tasks.
-(1) TUEV (event type classification) [harati2015improved]: An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
-The EEG signals are recorded at 23 channels and 256 Hz sampling rate.
-(2) TUAB (abnormal detection) [lopez2015automated]: A corpus of EEG recordings annotated as either normal or abnormal.
-Similar to TUEV, the EEG signals of TUAB are recorded at 23 channels and 256 Hz sampling rate.
-(3) SHU-MI (motor imagery classification) [ma2022large]: An EEG motor imagery dataset designed for decoding imagined movements.
-It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand.
-(4) CL-Drive (cognitive load assessment) [angkan2024multimodal]: It is a multimodal dataset designed for cognitive load assessment, which includes EEG signals.
-The EEG data are recorded using four channels at a sampling frequency of 256 Hz.
-This study explores the ternary classification of cognitive load into high, medium, and low levels.
-(5) Seed-VII (emotion recognition) [jiang2024seed]: It is a large-scale multimodal emotion dataset.
-Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise).
-The preprocessed EEG signals were downsampled to 200 Hz.
-(6) DMER (emotion distribution prediction) [yang2024multimodal]: It is a multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos).
-In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) [mackinnon1999short], reporting the intensity of ten emotional adjectives.
-Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution.
-The EEG recordings comprise 18 channels sampled at 300 Hz.
-It is worth noting that we primarily select downstream datasets from the past five years, a strategic shift from the benchmarks used in works like CBraMod [wang2024cbramod] or EEGPT [wangeegpt].
-This decision stems from the need to evaluate the model against the fidelity and complexity of modern neuroscience standards.
-Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms — ranging from naturalistic stimuli to multimodal retrieval — better assess the capacity for generalizable, high-level cognitive decoding.
-However, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets.
-Despite falling outside the five-year window, these two widely adopted benchmarks are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
-IV-B Experimental Setup
-For the pre-training dataset TUEG, following prior work [wang2024cbramod], we select 19 common EEG channels.
-All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels × 6,000 time points).
+Downstream tasks and Datasets.
+To comprehensively evaluate our method, we consider multiple downstream tasks, including event type classification (TUEV dataset (Harati et al., 2015)), abnormal detection (TUAB dataset (Lopez et al., 2015)), motor imagery classification (SHU-MI dataset (Ma et al., 2022)), cognitive load assessment (CL-Drive dataset (Angkan et al., 2024)), emotion recognition (Seed-VII dataset (Jiang et al., 2024b)), and emotion distribution prediction (DMER dataset (Yang et al., 2024b)).
+Detailed descriptions of these datasets are provided in the supplementary material.
+Notably, we primarily select downstream datasets from the past five years, representing a strategic shift from the benchmarks used in prior works such as CBraMod (Wang et al., 2025) and EEGPT (Wang et al., 2024).
+This choice is motivated by the need to evaluate the model against the fidelity and complexity of contemporary neuroscience standards.
+Older datasets often lack the high spatial density and signal quality required to stress-test a foundation model, whereas recent paradigms—ranging from naturalistic stimuli to multimodal retrieval—offer a more effective assessment of the model’s capacity for generalizable, high-level cognitive decoding.
+Nevertheless, to ensure fair comparison with established baselines, we make exceptions for the TUAB and TUEV datasets.
+Although these two widely adopted benchmarks fall outside the five-year window, they are retained as standardized anchors, allowing us to validate our model’s performance on classical clinical tasks alongside its capabilities on modern, complex scenarios.
+4.2. Experimental Setup
+Data Pre-processing.
+For the pre-training dataset TUEG, following prior work (Wang et al., 2025), we select 19 common EEG channels.
+All EEG signals are resampled to 200 Hz and segmented into non-overlapping 30-second samples (19 channels $\times$ 6,000 time points).
 Each EEG patch is set to a temporal length of 1 second, resulting in 570 patches per sample.
-For all downstream datasets, we maintain consistency with our pre-training data by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
-For the downstream datasets TUEV and TUAB, consistent with previous studies [wangeegpt, jiang2024large, wang2024cbramod], we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively.
-We adopt the train-test splits provided by these datasets.
-For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold subject-wise cross-validation.
-For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples.
-We employ a subject-independent split strategy for these datasets.
-We fine-tune EEGPT [wangeegpt], LaBraM-Base [jiang2024large] (the only variant with available pre-trained weights), and CBraMod [wang2024cbramod] based on their publicly available code and pre-trained weights, except when their experimental results are already reported in the original papers.
-For the EEG generator (diffusion module) and encoder, we adopt the DiT and ViT architectures, respectively.
-Following common practices in generative modeling literature, we maintain an exponential moving average (EMA) of model weights with a decay rate of 0.9999 during pre-training.
-The specific hyperparameters for pre-training are detailed in Table I.
-We evaluate the representation learning capabilities of EEGDM through fine-tuning on downstream tasks.
-Specifically, we predict corresponding labels using a linear predictor based on the outputs of the EEG encoder.
-Following the common evaluation protocols, we present results under the fine-tuning setting, where the predictor is jointly trained with the EEG encoder (i.e., the pre-trained encoder is not frozen).
-We employ cross-entropy loss for downstream classification tasks and KL divergence loss for downstream distribution prediction tasks.
-The hyperparameters for fine-tuning EEGDM on downstream tasks are provided in Table II.
-We conduct a series of quantitative and qualitative experiments to evaluate the proposed model, demonstrating its strong representation capabilities across diverse downstream datasets.
-TABLE I: Hyperparameters for EEGDM pre-training.
-Hyperparameter Value
-Optimization Batch size 32
-Peak learning rate 5e-4
-Minimal learning rate 5e-5
-Learning rate scheduler OneCycleLR
-Weight decay 0.05
-EMA decay rate 0.9999
-Gradient clipping norm 1
-Optimizer AdamW
-${\beta}_{1}$ 0.9
-${\beta}_{2}$ 0.999
-EEG Generator Layers N 8
-Hidden size 512
-Heads 8
-Normalization type LayerNorm
-EEG Encoder Layers 8
-Hidden size 512
-MLP size 2048
-Heads 8
-Sampling Classifier-free guidance 9
-Diffusion training steps 1000
-TABLE II: Hyperparameters for downstream fine-tuning. ($\star$) Depends on the dataset.
-Hyperparameter Value
-Batch size 64
-Peak learning rate 5e-4
-Minimal learning rate 5e-5
-Learning rate scheduler OneCycleLR
-Optimizer AdamW
-${\beta}_{1}$ 0.9
-${\beta}_{2}$ 0.999
-Weight decay 0.05
-Total epochs 30, 150
-Warmup Epochs 6, 30
-IV-C Results and Analysis
-In this section, we show that EEGDM achieves competitive performance across diverse downstream tasks, demonstrating the effectiveness of diffusion-based representation learning for EEG analysis.
-All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics.
-The reported model size corresponds to the downstream networks used for fine-tuning, as the generative module is not involved in downstream tasks.
-Results are compared against baselines, as shown in Tables III, IV, V and VI.
-Compared to other approaches, EEGDM consistently maintains either the best or second-best performance in key metrics (balanced accuracy or KL divergence).
-TABLE III: Performance comparison on TUEV. The best metric is highlighted in bold, and the second-best metric is underlined.
+For all downstream datasets, we maintain consistency with our pre-training setup by resampling EEG signals to 200 Hz and setting each EEG patch duration to 1 second.
+Table 1. Performance comparison on TUEV. The best result is highlighted in bold, and the second-best result is underlined.
 Methods Params Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
@@
 […]
-[…]
-TABLE IV: Performance comparison on TUAB. The best metric is highlighted in bold, and the second-best metric is underlined.
+Table 2.
+Performance comparison on TUAB.
+The best result is highlighted in bold, and the second-best result is underlined.
+NA indicates that the result is not available in the original literature (Wang et al., 2024).
 Methods Params Balanced Accuracy ↑ AUROC ↑ AUC-PR ↑
@@
 […]
-[…]
-[…]
-TABLE V: Performance comparison on DMER. The best metric is highlighted in bold, and the second-best metric is underlined.
+Baselines and Implementation.
+We compare our EEGDM against state-of-the-art self-supervised EEG representation learning frameworks, including EEGPT (Wang et al., 2024), LaBraM-Base (Jiang et al., 2024c) (the only variant with available pre-trained weights), CBraMod (Wang et al., 2025), and Puah-EEGDM-SSM (Puah et al., 2025).
+All baselines are fine-tuned using their publicly available code and pre-trained weights, except in cases where results were directly reported in the original papers.
+For our model, we adopt the DiT architecture for the EEG generator (diffusion module) and a ViT backbone for the encoder.
+The model is pre-trained on a cluster of four NVIDIA A100 GPUs with a batch size of 32.
+The entire pre-training phase encompasses 329,400 steps, totaling approximately 112.96 wall-clock hours.
+Evaluation Protocols.
+We evaluate our method under two settings: fine-tuning and linear probing (with a frozen encoder), to comprehensively assess the quality of the learned representations.
+Extensive experiments are conducted to evaluate the proposed model, demonstrating its strong representation capability across diverse downstream datasets.
+Detailed hyperparameter configurations, data splitting strategies, and implementation specifics for both pre-training and downstream evaluation protocols are provided in the supplementary material.
+Table 3. Performance comparison on DMER. The best result is highlighted in bold, and the second-best result is underlined.
 Methods Chebyshev ↓ Clark ↓ Canberra ↓ KL ↓ Cosine ↑ Intersection ↑
@@
 […]
-TABLE VI: The results of different methods on various datasets. The best metric is highlighted in bold, and the second-best metric is underlined.
+[…]
+[…]
+Table 4.
+Performance comparison on SHU-MI, CL-Drive and Seed-VII datasets.
+The best result is highlighted in bold, and the second-best result is underlined.
+Results for Puah-EEGDM-SSM on SHU-MI are omitted due to performance near chance levels.
 Datasets Methods Balanced Accuracy ↑ Weighted F1 / AUROC ↑ Cohen’s Kappa / AUC-PR ↑
@@
 […]
-TABLE VII: The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
+[…]
+[…]
+[…]
+[…]
+[…]
+4.3. Results and Analysis
+In this section, we demonstrate that EEGDM achieves competitive performance across diverse downstream tasks, underscoring the effectiveness of diffusion-based representation learning for EEG analysis.
+All experiments are conducted with multiple random seeds, and we report the mean and standard deviation of evaluation metrics.
+A comprehensive evaluation reveals that among all compared methods, our EEGDM is the only model that attains a top-two position on all six downstream datasets (except for EEGDM (T-Pre) on TUEV) with respect to the primary evaluation metrics (balanced accuracy and KL divergence).
+As shown in Table 1, EEGDM achieves a balanced accuracy of 68.19% on the TUEV dataset when pre-trained on the large-scale TUEG corpus, outperforming several state-of-the-art baselines.
+Notably, while the general-purpose version of EEGDM trails the results reported by Puah-EEGDM-SSM (Puah et al., 2025) (75.57%) on this specific benchmark, it is essential to consider the difference in pre-training data distributions.
+Unlike our model, which targets universal feature extraction across diverse domains (as evidenced by its superior results across other downstream datasets), Puah-EEGDM-SSM was pre-trained and evaluated exclusively on the TUEV dataset, benefiting from a high degree of domain alignment.
+To ensure a rigorous and fair architectural comparison, we introduced a task-specific variant of our model, denoted as EEGDM (T-Pre), which follows the same protocol as Puah-EEGDM-SSM (Puah et al., 2025) by utilizing TUEV for both pre-training and downstream evaluation.
+Under this aligned setting, EEGDM (T-Pre) achieves a state-of-the-art accuracy of 79.71%.
+This result demonstrates the superior representational capacity of our Latent Diffusion Transformer (DiT) architecture operating on a refined manifold.
+It further indicates that the performance gap in the general setting is a byproduct of domain shift in pre-training data rather than architectural limitations.
+The discrepancy between balanced accuracy and F1-score suggests that EEGDM’s latent space prioritizes holistic representation over majority-class optimization, leading to better pattern diversity at the cost of higher sensitivity to class boundary noise.
+Table 2 shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset.
+Securing the second-best performance on those metrics, EEGDM demonstrates competitive performance with a narrow margin.
+The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well to binary abnormality detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
+Table 3 presents the results for the DMER dataset, which uses distributional metrics.
+EEGDM achieves the best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second.
+The strong performance on DMER, a task requiring the prediction of an emotion distribution rather than a discrete class, strongly suggests that diffusion-based pre-training learns rich, fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in EEG signals.
+Table 4 further validates the generalizability and robustness of EEGDM across diverse BCI tasks.
+Specifically, on the SHU-MI dataset, EEGDM ranks second in balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod.
+Notably, the results for Puah-EEGDM-SSM (Puah et al., 2025) on this specific benchmark are omitted from the table, as its performance fluctuated near chance levels (e.g., accuracy $\approx 50\%$).
+This discrepancy suggests that while domain-specific pre-training (as used in (Puah et al., 2025)) may excel in aligned tasks, our Latent Diffusion Transformer (DiT) architecture, pre-trained on a broader corpus, exhibits superior cross-task adaptability, particularly in capturing the subtle spatial-temporal patterns required for motor imagery.
+On the CL-Drive dataset, EEGDM achieves the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities.
+On the Seed-VII dataset, EEGDM continues its strong performance with the best balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
+Table 5. Performance comparison of two data augmentation strategies on the CL-Drive dataset.
+Strategy Balanced Accuracy ↑ Weighted F1 ↑ Cohen’s Kappa ↑
+[…]
+[…]
+Beyond representation learning, a distinctive advantage of our proposed EEGDM is its generative capability to synthesize high-fidelity EEG data, an ability inherently absent in mask-reconstruction frameworks such as MAE.
+This ability is particularly effective in mitigating the pervasive issues of data sparsity and class imbalance in neural decoding.
+We evaluate this capability on the CL-Drive dataset—a relatively small-scale benchmark characterized by significant class skewness—using two augmentation strategies: isoportion expansion and balanced augmentation.
+As summarized in Table 5, our generative augmentation significantly improves downstream performance.
+By populating under-represented manifolds with synthetic yet semantically consistent samples, EEGDM effectively rectifies biased decision boundaries caused by data scarcity.
+These results underscore that our diffusion-based pre-training not only distills robust features but also serves as a powerful plug-and-play data engine, enhancing the model’s adaptability to real-world, long-tailed EEG distributions.
+Comprehensive qualitative evaluations, including t-SNE visualization and EEG signal reconstruction, are provided in the supplementary material to further validate the representation quality and generative fidelity of EEGDM.
+4.4. Ablation Study
+To verify the individual and collective contributions of PCA and channel augmentation, we evaluate four variants of EEGDM, as summarized in Table 6.
+Our analysis reveals a compelling synergistic relationship between these two components that extends beyond simple additive effects.
+Interestingly, the results indicate that employing either PCA or channel augmentation in isolation provides marginal improvements or even leads to slight performance degradation.
+While PCA projects signals into a latent space with improved signal-to-noise ratio (SNR), the reduction in dimensionality may inadvertently discard fine-grained temporal nuances.
+Without sufficient input diversity, the encoder tends to underfit the complex underlying brain states.
+Conversely, although channel augmentation enriches the training perspectives, it significantly increases the optimization burden when the model is tasked with reconstructing raw, noisy signals.
+The state-of-the-art performance achieved by Variant D underscores a powerful complementary effect.
+PCA simplifies the latent manifold, making the model more sensitive to the structural information provided by the augmented views.
+Meanwhile, the diverse augmented inputs prevent the model from losing critical information—effectively compensating for the potential information loss incurred by dimensionality reduction.
+To further assess the generalizability of this mechanism, we conducted a low-resource fine-tuning experiment on the TUAB dataset using only 10% of the training data (last column); we also use this setting in Fig.
+3.
+The results consistently show that the integration of PCA and EEG augmentation remains vital even in small-scale data settings.
+Table 6. The results of the ablation studies for PCA and EEG augmentation, using the metric of average balanced accuracy.
 Variants PCA EEG Augmentation TUEV TUAB (10%)
@@
 […]
-On the TUEV dataset, EEGDM achieves a balanced accuracy of 68.19%, outperforming both the previous state-of-the-art method CBraMod (66.71%) and other baselines.
-This superiority highlights the exceptional discriminative power of the representations learned by the diffusion model for classifying different event types in EEG signals.
-Table IV shows that EEGDM achieves a balanced accuracy of 81.80% and an AUC-PR of 90.85% on the TUAB dataset.
-Securing the second-best performance on those metrics, EEGDM overall demonstrates comparable performance with a competitive margin.
-The consistent performance across both balanced accuracy and AUC-PR metrics suggests that EEGDM learns robust representations that generalize well for binary abnormal detection tasks, maintaining high discriminative power even in challenging clinical scenarios.
-Table V presents the results for the DMER dataset, which uses distributional metrics.
-EEGDM achieves best performance across various divergence and similarity metrics, with the exception of the Clark metric, where it ranks second.
-The overall high performance on DMER, a task requiring the prediction of a continuous emotion distribution rather than a discrete class, strongly suggests that the diffusion-based pre-training learns rich and fine-grained features that preserve the complex, subtle, and continuous emotional states embedded in the EEG signals.
-Table VI further validates the generalizability and robustness of EEGDM across diverse BCI tasks.
-Specifically, on the SHU-MI dataset, EEGDM ranks second in terms of balanced accuracy (63.24%) and AUC-PR (70.70%), while achieving the best performance in AUROC (69.96%), demonstrating competitive performance compared with CBraMod.
-For the CL-Drive dataset, EEGDM shows the best balanced accuracy and the best Cohen’s Kappa, demonstrating superior cognitive load feature extraction capabilities.
-On the Seed-VII dataset, EEGDM continues its strong performance with the best performance in balanced accuracy and ranks second in weighted F1 and Cohen’s Kappa, further validating its effectiveness in emotion-related tasks.
-These comprehensive results across diverse datasets validate the effectiveness of using EEG signal generation as a self-supervised pre-training objective, demonstrating that the combination of encoder and diffusion denoiser, enhanced by PCA and channel augmentation, successfully learns high-quality EEG representations for improved downstream task performance.
 (a) TUEV
 (b) TUAB
-Figure 3: Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
-IV-D Ablation Study
-We conduct ablation studies to analyze the contributions of PCA and channel augmentation components, respectively, by setting up four variant versions as shown in Table VII, in which the average balanced accuracy on three random seeds is reported.
-Variant D consistently achieves the best performance across all settings, demonstrating that PCA and EEG augmentation are important for our model.
-However, we find that using PCA or EEG augmentation alone is either detrimental to or has minimal impact on our model’s performance.
-This may be attributed to the fact that PCA reduces the dimensionality of the source signal, making the model more sensitive to the diversity of conditional information, while data augmentation provides sufficient diversity to support learning in this reduced-dimensional space.
-Consequently, the PCA and EEG augmentation operations achieve complementary effects in terms of generalization performance and accuracy.
-Furthermore, for TUAB, we randomly selected 10% of the training set for fine-tuning and tested it on the entire test set (see the last column in Table VII); we also use this setting in Fig.
-3.
-The results show that the advantages of incorporating PCA and EEG augmentation remain evident even when small-scale datasets are used for fine-tuning.
+Figure 3. Performance comparison with different numbers of PCA components on TUEV and TUAB datasets.
 We also investigate the impact of the number of PCA components on downstream task performance.
 Fig.
-3 presents results on TUEV and TUAB with varying numbers of PCA components, which reveal a nonlinear relationship between the number of PCA components and downstream task performance.
-In particular, insufficient number of components may result in large information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance.
+3 presents results on TUEV and TUAB with varying numbers of PCA components, revealing a non-monotonic relationship between the number of PCA components and downstream task performance.
+Specifically, an insufficient number of components may result in substantial information loss, whereas an excessive number introduces noise contamination, both of which degrade model performance.
 To balance these trade-offs, we empirically select an intermediate value of 20 components, which yields optimal performance in our experiments.
-(a) TUEV
-(b) TUAB
-Figure 4: The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
-Figure 5: Visualization of reconstructed EEG signals and their PSD.
-IV-E Visualization
-To evaluate the quality of learned representations, we visualize the EEG data embeddings using t-SNE [van2008visualizing] as presented in Fig.
+In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG generation, as validated by its effectiveness in augmenting an imbalanced and small-scale dataset; (2) Effective representation learning, as demonstrated by competitive linear probing results and substantial performance gains through fine-tuning; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent improvements in balanced accuracy and robustness over state-of-the-art baselines.
+5. Conclusion
+We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation.
+Our approach adopts the generative task as the self-supervised EEG representation learning objective.
+By (1) integrating a denoiser with the EEG encoder and (2) leveraging PCA-based projection and EEG channel augmentation, EEGDM performs the generation task in the latent space using rich conditional information while simultaneously learning powerful EEG semantic representations that can be applied to diverse downstream tasks.
+Comprehensive experimental results demonstrate the model’s effectiveness across multiple downstream tasks.
+Supplementary Material
+Appendix A Dataset Details for Downstream Tasks
+This section provides a detailed description of the downstream datasets employed to evaluate the performance of our proposed model.
+(1) TUEV (event type classification) (Harati et al., 2015): An EEG dataset containing six event categories: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
+The EEG signals are recorded with 23 channels at a sampling rate of 256 Hz.
+(2) TUAB (abnormal detection) (Lopez et al., 2015): A corpus of EEG recordings annotated as either normal or abnormal.
+Similar to TUEV, the EEG signals in TUAB are recorded with 23 channels at a sampling rate of 256 Hz.
+(3) SHU-MI (motor imagery classification) (Ma et al., 2022): An EEG motor imagery dataset designed for decoding imagined movements.
+It features 32 EEG channels sampled at 250 Hz, including two types of motor imagery tasks: left hand and right hand.
+(4) CL-Drive (cognitive load assessment) (Angkan et al., 2024): A multimodal dataset designed for cognitive load assessment, which includes EEG signals.
+The EEG data are recorded using four channels at a sampling frequency of 256 Hz.
+This study explores the ternary classification of cognitive load into high, medium, and low levels.
+(5) Seed-VII (emotion recognition) (Jiang et al., 2024b): A large-scale multimodal emotion dataset.
+Seed-VII includes 62-channel EEG and eye-movement data from 20 healthy subjects across seven emotion categories (disgust, fear, sad, neutral, happy, anger, surprise).
+The preprocessed EEG signals were downsampled to 200 Hz.
+(6) DMER (emotion distribution prediction) (Yang et al., 2024b): A multimodal emotion recognition corpus that includes EEG signals along with other modalities (GSR, PPG, and frontal facial videos).
+In each trial, participants watched video clips and subsequently completed a self-assessment using a ten-item version of the Positive Affect and Negative Affect Schedule (PANAS) (Mackinnon et al., 1999), reporting the intensity of ten emotional adjectives.
+Each emotion rating ranges from 1 (not at all) to 5 (extremely), and the resulting scores are transformed into a 10-dimensional emotion distribution.
+The EEG recordings comprise 18 channels sampled at 300 Hz.
+Appendix B Detailed Experimental Settings for Downstream Datasets
+For the downstream datasets TUEV and TUAB, consistent with previous studies (Wang et al., 2024; Jiang et al., 2024c; Wang et al., 2025), we utilize 16 common bipolar montage channels from the international 10-20 system, with sample durations of 5 seconds and 10 seconds, respectively.
+We adopt the train-test splits provided by these datasets.
+For the CL-Drive dataset, we use the channel configuration and sample duration (10 seconds) provided by the dataset and apply 10-fold cross-validation.
+For the SHU-MI, DMER, and SEED-VII datasets, we use the channel configurations provided by each dataset and segment the signals into 4-second samples.
+A subject-independent split strategy is adopted for these datasets.
+Appendix C Implementation Details
+Table 7. Hyperparameters for EEGDM pre-training.
+Hyperparameter Value
+Optimization Batch size 32
+Peak learning rate 5e-4
+Minimal learning rate 5e-5
+Learning rate scheduler OneCycleLR
+Weight decay 0.05
+EMA decay rate 0.9999
+Gradient clipping norm 1
+Optimizer AdamW
+$\beta$} (0.9, 0.999)
+Diffusion Module Layers N 8
+Hidden size 512
+Heads 8
+Normalization type LayerNorm
+EEG Encoder Layers 8
+Hidden size 512
+MLP size 2048
+Heads 8
+Sampling Classifier-free guidance 9
+Diffusion training steps 1000
+Pre-training: The pre-training of EEGDM is conducted on the large-scale TUEG dataset to learn robust, general-purpose neural representations.
+We utilize the AdamW optimizer with a peak learning rate of 5e-4, which is decayed to a minimal value of 5e-5 following a OneCycleLR scheduling policy.
+We apply an Exponential Moving Average (EMA) on the model weights with a decay rate of 0.9999.
+The architecture consists of an 8-layer EEG Encoder and an 8-layer Diffusion Module, both featuring a hidden size of 512 and 8 attention heads.
+Detailed hyperparameter configurations for the optimization process, model architecture, and diffusion sampling are summarized in Table 7.
+Downstream tasks: We evaluate the pre-trained EEGDM across diverse downstream scenarios using two protocols: fine-tuning and linear probing.
+In the fine-tuning setting, both the pre-trained encoder and the task-specific prediction head are jointly optimized.
+For linear probing, the EEG encoder remains strictly frozen, and only the prediction head is trained to assess the quality of the learned representations.
+To accommodate the varying scales and complexities of different downstream datasets, we employ task-specific hyperparameter configurations, as detailed in Table 8.
+All downstream experiments are conducted using the AdamW optimizer to ensure training stability.
+Detailed hyperparameters for both protocols, including batch sizes, learning rates, and weight decay, are summarized in Table 8.
+Table 8. Hyperparameters for downstream fine-tuning and linear probing. ($\star$) Depends on the dataset.
+Hyperparameter Fine-tuning Linear Probing
+Encoder status Unfrozen Frozen
+Learning rate 5e-4 2e-4, 5e-4, 2e-3
+Total epochs 30, 150 30, 50, 100
+Weight decay 0.05 0.05, 0.02, 0.01
+Batch size 64 32, 64
+Optimizer AdamW AdamW
+Learning rate scheduler OneCycleLR OneCycleLR, ReduceLROnPlateau, CosineAnnealingLR
+$\beta$} (0.9, 0.999) (0.9, 0.999)
+Warmup epochs 6, 30 3, 10
+Appendix D Qualitative Analysis
+To evaluate the quality of the learned representations, we visualize the EEG data embeddings using t-SNE (Van der Maaten and Hinton, 2008) as presented in Fig.
 4.
@@
 This clear clustering structure further demonstrates the EEGDM’s ability to learn generic yet informative EEG representations that can be successfully adapted to diverse downstream tasks.
-Beyond representation learning, EEGDM also achieves high-quality EEG signal reconstruction, as evidenced by visualizations of reconstructed signals and their power spectral density (PSD) analyses.
+To further validate the generative quality, we provide visualizations of reconstructed signals and their Power Spectral Density (PSD).
 As shown in Fig.
@@
 These results indicate that the proposed model can generate realistic EEG signals and effectively capture high-level semantic information from EEG data.
-In summary, our experimental results highlight three key advantages of EEGDM: (1) High-quality EEG reconstruction, as validated by signal reconstruction visualization; (2) Effective representation learning, as demonstrated by distinct class separation in t-SNE visualization; and (3) Competitive performance across diverse downstream tasks, as evidenced by consistent performance improvements across tasks.
-V Conclusion
-We present EEGDM, a novel latent diffusion framework for EEG representation learning, through explicit conditioning on EEG channel augmentation.
-Our EEGDM takes the generative task as the self-supervised EEG representation learning objective.
-By (1) integrating a denoiser with the EEG encoder and (2) making use of PCA-based projection and EEG channel augmentation, EEGDM can perform the generation task in the latent space using rich conditional information while learning powerful EEG semantic representations that can be applied to diverse downstream tasks.
-Comprehensive experimental results demonstrate the model’s effectiveness across multiple downstream tasks.
+(a) TUEV
+(b) TUAB
+Figure 4. The t-SNE visualizations. Clear inter-class separation and intra-class compactness indicate that the learned representations are discriminative.
+Figure 5. Visualization of reconstructed EEG signals and their PSD.